Suppose S,TL(V) are self-adjoint. Prove that ST is self-adjoint if and only if ST=TS.
1
Expert's answer
2021-07-27T16:02:27-0400

Given S,T "\isin" L(V) are self-adjoint.

If suppose ST = TS,

"\implies" ST + TS = 2(ST). Since 2(ST) is self-adjoint, where 2 is a real number.

consider, 2"\langle" ST(v), w "\rangle" = "\langle" 2ST(v), w "\rangle" = "\langle" v, 2ST(w) "\rangle" = 2"\langle" v, ST(w) "\rangle"

therefore, we get,

"\implies" "\langle" ST(v), w "\rangle" = "\langle" v, ST(w) "\rangle" "\implies"  ST is self-adjoint. 


Now suppose ST is self-adjoint. then "\langle" ST(v), w "\rangle" = "\langle" v, ST(w) "\rangle"

and "\langle" ST(v), w "\rangle" = "\langle" v,(ST)*w "\rangle" = "\langle" v, T*S*w "\rangle" = "\langle" v, TS(w) "\rangle" (since T, S are self-adjoint)

we get, "\langle" v, ST(w) "\rangle" = "\langle" v, TS(w) "\rangle" for all  v, w ∈ V

"\implies" "\langle" v, (ST-TS)(w) "\rangle" = 0  for all v, w ∈ V ,

so setting v = (ST − T S)w

"\implies" "\langle" (ST-TS)(w) , (ST-TS)(w) "\rangle" = 0

"\implies" ||(ST − T S)w||2 = 0 for all w ∈ V ,

therefore,

ST − T S = 0  "\implies" ST = TS